An ore of Fe has formula Fe0.93O1.0. The mole fraction of total iron present in the form fe (ii) is ?
(o.93 and 1.0 are subscripts )

Grade:11

1 Answers

Beenu Mathew
7 years ago
assume 100 moles of wustite; therefore you have 93 moles of Fe (Fe3+ plus Fe2+) and 100 moles of oxygen 2- 

x= moles of Fe3+ 
y= moles of Fe2+ 
z= 100 moles of O; oxygen has a charge of 2- 
now,
x+y=93   
y=93-x
NOW WE ALL KNOW THAT COMPOUNDS ARE SUPPOSED TO BE ELECTRICALLY NEUTRAL
x(3) + 2(93-x) + -2(z)=0
x(3) + 2(93-x) = 2(z)
z=100
x=14 (on solving)
y=93-14 = 79
mole fraction of total iron present in the form of Fe(ii)
is 79/14+79 = .84
PLS APPROVE
HOPE THIS HELPS
THANK YOU
BEST OF LUCK
YAHWeG